OpenStudy (anonymous):

Radical equations. -8+(sqrt of 5a-5)=-3

OpenStudy (anonymous):

\[-8+\sqrt{5a-5}=-3\] Add 8 to both sides \[\sqrt{5a-5}=5\] Square bot sides 5a-5=25 add 5 to both sides 5a=30 Divide by 5 a=30/5 a=6

OpenStudy (anonymous):

@moneylove01

OpenStudy (anonymous):

@31356

OpenStudy (31356):

Correct! :D